OpenStudy (anonymous):

\(\ \Huge \ln{5-2x} = -3? \)

OpenStudy (anonymous):

That's ln(5-2x)

OpenStudy (kropot72):

\[\ln (5-2x)=-3\] \[e ^{-3}=5-2x\] \[2x=5-\frac{1}{e ^{3}}\]